OpenStudy (anonymous):

Expand each logarithm log3squareroot(2x) please help

OpenStudy (ash2326):

\[\log_3 \sqrt{2x}\] so it's \[\large \log_3 (2x)^{\frac{1}{2}}\] We know the logarithm property \[\log_a b^c=c \times \log_a b\] Using this here \[\frac{1}{2}\large \log_3 2x\] Do you get this?

OpenStudy (anonymous):

not really idk how to get the anwser

OpenStudy (ash2326):

Do you know logarithms and their properties ?

OpenStudy (anonymous):

not well

OpenStudy (ash2326):

Logarithms are defined as \[\large \log_x y=z\] This implies \[\large y=x^z\] Some important properties of logarithms \[1)\ \log a^b= b\times \log a\] \[2)\ \log ab=\log a+\log b\] \[3)\ \log \frac ab=\log a-\log b\] Base can be any number > 0 but not 1

OpenStudy (anonymous):

well can you help me figure out the anwser?

OpenStudy (ash2326):

@asdfqwer1234 I gotta go now, my friend will help you here @apoorvk please help here

OpenStudy (apoorvk):

Hey asdfqwer1234 - did you follow what Ash said above? If yes, good. If no, read it up very slowly, calmly and carefully. :)

OpenStudy (apoorvk):

The answer is actually staring at you :)

OpenStudy (anonymous):

what i don't get it?

OpenStudy (apoorvk):

okay - \[(2x)^{\frac 1 2} = \sqrt{2x} = \sqrt2 \times \sqrt x\] do you agree with this much?

OpenStudy (anonymous):

yeah

OpenStudy (apoorvk):

so can you use the second property that Ash posted above - the one that says- "log(a x b) = log(a) + log(b)"? Here, your 'a' would be sqrt(2) and 'b' sqrt(x).

OpenStudy (anonymous):

thanks

OpenStudy (apoorvk):

You sure you'll be able to do this stuff now?

OpenStudy (anonymous):

no not really but i think i will figure it out

OpenStudy (apoorvk):

Well, just tag me in or anyone online (am going off right now) if you need more help!